1) Solve the triangle: c=10, b=14, A=120 degrees
2) Take the area of the triangle a=12, b=12 and the included angle C= 60 degrees
3) Calculate the arc length of a sector of a circle with a radius of 7 and central angle measure of 98 degrees. Now, calculate the area of the sector.
Thank you!

Answer by stanbon(75887)

1) Solve the triangle: c=10, b=14, A=120 degrees
a^2 = b^2+c^2-2bc*cos(A)
a = sqrt[14^2+10^2-2*140(-1/2)]
a = sqrt[436] = 20.88
-----------------------------------
2) Take the area of the triangle a=12, b=12 and the included angle C= 60 degrees
Area = (1/2)a*b*cos(60) = (1/2)140*(1/2) = 35 sq units
------------------------------------
3) Calculate the arc length of a sector of a circle with a radius of 7 and central angle measure of 98 degrees. Now, calculate the area of the sector.
98 degrees = 98(pi/18) = 1.71 radians
arc length = 1.71*7 = 11.97
-----------------------
Area of circle = pi*7^2 = 49pi sq units
------
Area of sector = (98/360)49pi = 41.91 sq units
----------------------------
Cheers,
Stan H.
========================